# Towards Real-World Blind Face Restoration with Generative Diffusion Prior

### 25 Dec 2023 v1

@@ IV-A1 Implementation
Stable Diffusion 2.1-base is adopted as our foundational denoising networks and generative facial prior. We first finetune the frozen diffusion model for 100K iterations. Subsequently, we unfreeze the decoder weights of the U-Net in Stable Diffusion and train the whole restoration model for 150K iterations. The adamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$, weight decay = 0.01) is employed with the cosine annealing strategy, where the learning rate gradually decreases from the initial learning rate $1\times 10^{-4}$ to zero for the last 50K iterations. We set the batch size to 64 on 4 NVIDIA A100 GPUs in the training stage and apply the DDIM [65] sampler with 50 steps in the inference stage.
@@ IV-A2 Datasets
Similar to [11, 4, 3], we employ a synthetic paired dataset CelebA-Test and three real-world datasets: LFW-Test, CelebAdult-Test, and WebPhoto-Test to evaluate our BFRffusion. CelebA-Test consists of 3K paired images that are synthesized from CelebA-HQ [47] images. LFW-Test comprises 1,711 low-quality images which are the first images for each identity in the validation partition of LFW. CelebAdult-Test which consists of 180 adult faces, and WebPhoto-Test which consists of 407 real-life faces are all collected from the Internet. All these testing datasets are public and do not overlap with the training dataset.
@@ IV-B Comparison with SOTA Face Restoration Methods
TABLE II: Quantitative comparison on CelebA-Test for blind face restoration.Red and Blue indicate the best and the second-best performance.
@@
We compare the performance of our BFRffusion with several state-of-the-art face restoration methods: HiFaceGAN [69], DFDNet [1], PSFRGAN [9], GFP-GAN [11], GPEN [12], VQFR [4], CodeFormer [5]. Specifically, HiFaceGAN [69] formulates the face restoration task as a generative problem guided by semantics, and the problem is tackled by a multi-stage framework consisting of collaborative suppression and replenishment. DFDNet [1] generates a deep dictionary of key facial components extracted from high-quality images as the reference prior and selects similar features to degraded inputs to guide the restoration process. PSFRGAN [9] is a multi-scale progressive restoration network, that leverages both the geometric prior (parsing maps) and pixel space information (input degraded images). GFP-GAN [11] employs the pretrained StyleGAN2 [42], which encapsulates rich and diverse generative facial priors, and spatial feature transform layers to restore realistic and faithful faces. GPEN [12] is a two-stage framework, which first embeds learned GAN blocks into a U-shaped DNN as a prior decoder, and then fine-tunes the GAN prior with the synthesized low-quality face images. VQFR [4] is a VQ-based face restoration method equipped with the VQ codebook as a facial detail dictionary and the parallel decoder design. CodeFormer [5] is a Transformer-based prediction network, that models both the global composition and the context of low-quality faces for code prediction. It has shown superior robustness to degradation with the codebook prior. In summary, our selection of the above seven methods is guided by the three following criteria. First, these methods achieve state-of-the-art performance in terms of widely used metrics such as PSNR, SSIM, etc. Second, the testing source codes of these methods are publicly accessible. Third, these methods are specifically designed for blind face restoration.
TABLE III: Quantitative comparison on three real-world datasets for blind face restoration. Red and blue indicate the best and the second-best performance.
Dataset LFW-Test CelebAdult-Test WebPhoto-Test
@@
Input 137.56 118.33 170.11
HiFaceGAN [69] 64.50 104.02 116.12
DFDNet [1] 62.57 105.71 100.68
PSFRGAN [9] 51.89 104.08 88.45
GFP-GAN [11] 49.96 104.75 87.35
GPEN [12] 57.00 107.53 101.3
VQFR [4] 50.64 104.99 75.38
CodeFormer [5] 52.36 110.09 83.19
BFRffusion (ours) 49.92 103.64 84.29
@@ IV-B1 Comparisons on Synthetic Dataset
The quantitative comparisons of the methods mentioned above are reported in Table II. The results show that our BFRffusion achieves state-of-the-art performance on the CelebA-Test. Specifically, our BFRffusion achieves the best performance regarding pixel-wise metrics PSNR and SSIM. Furthermore, BFRffusion obtains the lowest Deg. and LDM scores, indicating its ability to accurately recover identity and facial details. In addition, BFRffusion achieves comparable LPIPS and FID scores, suggesting that the quality of restored faces is close to the ground truth. The qualitative results are presented in Fig. 5. Our BFRffusion leverages the proposed multi-level feature extraction module to extract image features during the restoration process, allowing it to capture clear details of the entire image. In contrast, the methods, that require face detection (DFDNet [1]) or parsing maps (PSFRGAN [9]), are difficult to restore faithful details in the whole image. Moreover, thanks to the remarkable generative ability of the pretrained Stable Diffusion, our BFRffusion successfully recovers realistic details in the eyes, mouth, decorations, etc. On the contrary, GAN-based methods [11, 12, 4, 5] are difficult to restore complex components, due to their limited generative ability. In conclusion, our BFRffusion performs better in the realness and fidelity of the facial details.
@@ IV-B2 Comparisons on Real-world Datasets
Furthermore, we compare the performance of the methods mentioned above in real-world scenarios based on three real-world datasets to evaluate the generalization ability. The quantitative comparisons are shown in Table III. Our BFRffusion achieves the best performance on both the LFW-Test and CelebAdult-Test datasets, while also demonstrating comparable performance on the WebPhoto-Test. The qualitative results are presented in Fig. 6. Our BFRffusion equipped with powerful generative prior is able to generate more realistic and detailed features such as hair, teeth, eyes, makeup, and decorations, amongst others. Although VQFR [4] obtains the lowest FID scores on the WebPhoto-Test, it could not generate realistic details (e.g., teeth, earrings) as shown in WebPhoto-Test in Fig. 6.
IV-C Ablation Studies
In this section, we analyze and discuss the effectiveness of the internal modules in our BFRffusion and the strategies employed during our training process. All models are trained on FFHQ [19] and tested on CelebA-Test. The quantitative results are shown in Table IV and the qualitative results are presented in Fig. 7.
TABLE IV: Ablation Studies of our BFRffusion
@@ IV-C1 Shallow Degradation Removal Module
IV-C2 Multi-level Feature Extraction Module
Our multi-level feature extraction module (MFEM) is constructed using transformer blocks with time conditions. To test the importance of the transformer blocks, we replaced them with time-embedded ResBlocks [51, 52], which caused a significant drop in performance. It suggests that the proposed transformer blocks are important for the extraction of global information and local information. Then we remove the time conditions of the transformer blocks. This also causes a performance drop, demonstrating that time information is necessary for the feature extraction and denoising process.
@@ IV-C3 Trainable Time-aware Prompt Module
The trainable time-aware prompt module (TTPM) consists of a trainable prompt and time embedding. Firstly, we remove the time conditions, and there is only a trainable prompt left. Then we replace TTPM with the frozen CLIP text encoder and set the input prompt to blank, which is equivalent to a fixed prompt and is similar to previous works [17, 18]. Table IV(e) shows that our TTPM with the time conditions yields 0.17 dB PSNR gain over only the trainable prompt. What’s more, Table IV(f) shows that the trainable prompt yields 0.11 dB PSNR gain over the fixed prompt.
@@ IV-C4 Pretrained Denoising U-Net Module
We adopt the pretrained denoising U-Net module (PDUM) of Stable Diffusion as the foundation denoising networks and generative facial prior. To demonstrate the effectiveness of the pretrained Stable Diffusion weights, we train our BFRffusion without them for initializing the U-Net denoiser. A significant drop is observed and it can’t generate realistic glasses as shown in Fig. 7(g), which shows that the generative ability from the pretrained Stable Diffusion is quite important for the restoration process. We employ a training strategy where we finetune the frozen diffusion model for 100K iterations, then unfreeze the decoder weights of the U-Net in Stable Diffusion and train the whole restoration model for 150K iterations. This training strategy allows our BFRffusion to achieve wonderful realness and fidelity while retaining the generation ability of Stable Diffusion. We freeze the whole U-Net all the time leading to a bad optimization of the restoration model and a slight performance drop. On the other hand, unfreezing U-Net results in the forgetting of prior knowledge of the pretrained Stable Diffusion and a noticeable performance drop. In the U-Net architecture of diffusion models, the encoder is responsible for gradually downsampling the input into high-dimensional feature representation and extracting rich semantic information, while the decoder gradually restores the spatial details of the image through upsampling operations and performs prediction. We use the proposed MFEM to extract features from the input low-quality images, so We choose to unfreeze the decoder rather than the encoder to achieve better restoration results. Table IV(j) shows that the unfrozen decoder yields 0.70 dB PSNR gain over the unfrozen encoder.
TABLE V: Quantitative comparison on CelebA-Test for blind face restoration trained on FFHQ and PFHQ dataset. Bold indicates better performance.
@@
HiFaceGAN FFHQ 0.4770 42.18 3.16 24.92 0.6195
PFHQ 0.3997 33.77 2.20 25.03 0.6547
PSFRGAN FFHQ 0.4240 39.69 3.41 24.71 0.6557
PFHQ 0.4278 36.13 2.52 25.27 0.6690
BFRffusion FFHQ 0.3621 30.91 1.99 26.20 0.6926
PFHQ 0.3602 31.03 1.98 26.13 0.6951
TABLE VI: Quantitative comparison on three real-world datasets for blind face restoration trained on FFHQ and PFHQ dataset. Bold indicates better performance.
Methods Dataset LFW-Test CelebAdult-Test WebPhoto-Test
@@
HiFaceGAN FFHQ 64.50 104.02 116.12
PFHQ 48.19 119.27 76.10
PSFRGAN FFHQ 51.89 104.08 88.45
PFHQ 51.83 131.84 76.15
BFRffusion FFHQ 49.92 103.64 84.29
PFHQ 50.07 103.45 77.92
Figure 8: Visual comparison results of methods trained on FFHQ dataset and PFHQ dataset. The first column is the results on the FFHQ dataset and the second column is the results on the PFHQ dataset of each method. Methods trained on the PFHQ dataset could produce more faithful details because of the balanced character.
IV-D Utility Evaluation of the Built Privacy-preserving Face Dataset
To evaluate our privacy-preserving face dataset PFHQ, we train three representative blind face restoration methods: HiFaceGAN [69] (non-prior-based methods), PSFRGAN [9] (geometric prior-based methods), and BFRffusion (generative prior-based methods) on PFHQ dataset and compare the performance of these methods trained on FFHQ dataset [19]. We test these methods on CelebA-Test and three real-world testing datasets mentioned in Section IV-B. The quantitative comparisons on the public synthetic dataset CelebA-Test are reported in Table V. Bold values indicate better performance. Specifically, HiFaceGAN [69] trained on our PFHQ dataset achieves better performance on five metrics than HiFaceGAN trained on our FFHQ dataset. PSFRGAN [9] trained on our PFHQ dataset achieves better performance regarding PSNR, SSIM, Deg., and LMD than it trained on our FFHQ dataset. BFRffusion trained on our PFHQ dataset achieves comparable performance to that when trained on our FFHQ dataset. The quantitative comparisons on the public real-world testing datasets are reported in Table VI. Three blind face restoration methods trained on our PFHQ dataset also achieve comparable performance to those trained on the FFHQ dataset. The qualitative results are presented in Fig. 8. Due to the balanced characteristics of our PFHQ dataset, methods trained on it could produce more faithful details on various population groups.
@@ V Conclusion and Future Work
In this work, we propose BFRffusion with delicately designed architecture to leverage amazing generative priors encapsulated in pretrained Stable Diffusion for blind face restoration. Our BFRffusion effectively restores realistic and faithful facial details and achieves state-of-the-art performance on both synthetic and real-world public testing datasets. Furthermore, we build a privacy-preserving paired face dataset called PFHQ with balanced race, gender, and age. Extensive experiments show that our PFHQ dataset can serve as an alternative to real-face datasets for training blind face restoration methods.

### 18 Mar 2024 v2

@@ IV-A1 Implementation
Stable Diffusion 2.1-base is adopted as our foundational denoising networks and generative facial prior. We employ a new training strategy compared to the strategy used in StableSR [15] and DiffBIR [16], which maintain the U-Net frozen throughout all training phases. We first finetune the frozen diffusion model for 100K iterations. Subsequently, we unfreeze the decoder weights of the U-Net in Stable Diffusion and train the whole restoration model for 150K iterations. The adamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$, weight decay = 0.01) is employed with the cosine annealing strategy, where the learning rate gradually decreases from the initial learning rate $1\times 10^{-4}$ to zero for the last 50K iterations. We set the batch size to 64 on 4 NVIDIA A100 GPUs in the training stage and apply the DDIM [71] sampler with 50 steps in the inference stage.
@@ IV-A2 Datasets
Similar to [9, 17, 3, 18, 16], we employ a synthetic paired dataset CelebA-Test and three real-world datasets: LFW-Test, CelebAdult-Test, and WIDER-Test to evaluate our BFRffusion. CelebA-Test consists of 3K paired images that are synthesized from CelebA-HQ [53] images. LFW-Test comprises 1,711 low-quality images which are the first images for each identity in the validation partition of LFW. CelebAdult-Test which consists of 180 adult faces is collected from the Internet. WIDER-Test consists of 970 severely degraded face images from the WIDER Face dataset [72]. All these testing datasets are public and do not overlap with the training dataset.
@@ IV-B Comparison with SOTA Face Restoration Methods
TABLE III: Quantitative comparison on CelebA-Test for blind face restoration. Red and Blue indicate the best and the second-best performance.
@@
We compare the performance of our BFRffusion with several state-of-the-art face restoration methods: HiFaceGAN [76], DFDNet [1], PSFRGAN [7], GFP-GAN [9], GPEN [10], VQFR [17], CodeFormer [18], DR2 [50], StableSR [15] and DiffBIR [16]. Specifically, HiFaceGAN [76] formulates the face restoration task as a generative problem guided by semantics, and the problem is tackled by a multi-stage framework consisting of collaborative suppression and replenishment. DFDNet [1] generates a deep dictionary of key facial components extracted from high-quality images as the reference prior and selects similar features to degraded inputs to guide the restoration process. PSFRGAN [7] is a multi-scale progressive restoration network, that leverages both the geometric prior (parsing maps) and pixel space information (input degraded images). GFP-GAN [9] employs the pretrained StyleGAN2 [48], which encapsulates rich and diverse generative facial priors, and spatial feature transform layers to restore realistic and faithful faces. GPEN [10] is a two-stage framework, which first embeds learned GAN blocks into a U-shaped DNN as a prior decoder, and then fine-tunes the GAN prior with the synthesized low-quality face images. VQFR [17] is a VQ-based face restoration method equipped with the VQ codebook as a facial detail dictionary and the parallel decoder design. CodeFormer [18] is a Transformer-based prediction network, that models both the global composition and the context of low-quality faces for code prediction. It has shown superior robustness to degradation with the codebook prior. DR2 [50] captures semantic information through iterative denoising steps and is robust against common degradation. StableSR [15] is a blind super-resolution method leveraging prior knowledge encapsulated in pretrained Stable Diffusion. DiffBIR [16] is a two-stage framework with the pretrained Stable Diffusion. In summary, our selection of the above methods is guided by the three following criteria. First, these methods achieve state-of-the-art performance in terms of widely used metrics such as PSNR, SSIM, etc. Second, the testing source codes of these methods are publicly accessible. Third, these methods are specifically designed for blind face restoration.
TABLE IV: Quantitative comparison on three real-world datasets for blind face restoration. Red and blue indicate the best and the second-best performance.
Dataset LFW-Test CelebAdult-Test WIDER-Test
@@
Input 137.56 118.33 202.06
HiFaceGAN [76] 64.50 104.02 125.56
DFDNet [1] 62.57 105.71 57.84
PSFRGAN [7] 51.89 104.08 51.16
GFP-GAN [9] 49.96 104.75 40.59
GPEN [10] 57.00 107.53 46.99
VQFR [17] 50.64 104.99 37.87
CodeFormer [18] 52.36 110.09 39.06
DR2 [50] 49.85 107.70 54.34
StableSR [15] 41.15 105.09 33.05
DiffBIR [16] 39.58 105.02 32.35
BFRffusion (ours) 49.92 103.64 56.97
@@ IV-B1 Comparisons on Synthetic Dataset
The quantitative comparisons of the methods mentioned above are reported in Table III. The results show that our BFRffusion achieves state-of-the-art performance on the CelebA-Test. Specifically, our BFRffusion achieves the best performance regarding pixel-wise metrics PSNR and SSIM. Furthermore, BFRffusion obtains the lowest Deg. and LDM scores, indicating its ability to accurately recover identity and facial details. In addition, BFRffusion achieves comparable LPIPS and FID scores, suggesting that the quality of restored faces is close to the ground truth. The qualitative results are presented in Fig. 6. Our BFRffusion leverages the proposed multi-scale feature extraction module to extract image features during the restoration process, allowing it to capture clear details of the entire image. In contrast, the methods, that require face detection (DFDNet [1]) or parsing maps (PSFRGAN [7]), are difficult to restore faithful details in the whole image. Moreover, thanks to the remarkable generative ability of the pretrained Stable Diffusion, our BFRffusion successfully recovers realistic details in the eyes, mouth, decorations, etc. On the contrary, GAN-based methods [9, 10, 17, 18] are difficult to restore complex components, due to their limited generative ability. In conclusion, our BFRffusion performs better in the realness and fidelity of the facial details.
@@ IV-B2 Comparisons on Real-world Datasets
Furthermore, we compare the performance of the methods mentioned above in real-world scenarios based on three real-world datasets to evaluate the generalization ability. The quantitative comparisons are shown in Table IV. Our BFRffusion achieves the best performance on the CelebAdult-Test datasets, while also demonstrating comparable performance on the LFW-Test. However, our BFRffusion achieves poor performance on the WIDER-Test with severe degradation. We use up to 8x downsampling to simulate low-resolution situations and our model struggles to restore such highly degraded images in WIDER-Test. The qualitative results are presented in Fig. 7. Our BFRffusion equipped with powerful generative prior is able to generate more realistic and detailed features such as teeth, eyes, makeup, and decorations. Although DiffBIR [16] obtains the lowest FID scores on the LFW-Test and WIDER-Test, it notably lacks fine facial texture details and tends to appear smoother as shown in Fig. 7.
TABLE V: Ablation Studies of our BFRffusion
@@
IV-C Ablation Studies
In this section, we analyze and discuss the effectiveness of the internal modules in our BFRffusion and the strategies employed during our training process. All models are trained on FFHQ [19] and tested on CelebA-Test. The quantitative results are shown in Table V and the qualitative results are presented in Fig. 8.
@@ IV-C1 Shallow Degradation Removal Module
IV-C2 multi-scale Feature Extraction Module
Our multi-scale feature extraction module (MFEM) is constructed using transformer blocks with time conditions. To test the importance of the transformer blocks, we replaced them with time-embedded ResBlocks [57, 58], which caused a significant drop in performance. It suggests that the proposed transformer blocks are important for the extraction of global information and local information. Then we remove the time conditions of the transformer blocks. This also causes a performance drop, demonstrating that time information is necessary for the feature extraction and denoising process.
@@ IV-C3 Trainable Time-aware Prompt Module
The trainable time-aware prompt module (TTPM) consists of a trainable prompt and time embedding. Firstly, we remove the time conditions, and there is only a trainable prompt left. Then we replace TTPM with the frozen CLIP text encoder and set the input prompt to blank, which is equivalent to a fixed prompt and is similar to previous works [15, 16]. Table V(e) shows that our TTPM with the time conditions yields 0.17 dB PSNR gain over only the trainable prompt. What’s more, Table V(f) shows that the trainable prompt yields 0.11 dB PSNR gain over the fixed prompt.
@@ IV-C4 Pretrained Denoising U-Net Module
We adopt the pretrained denoising U-Net module (PDUM) of Stable Diffusion as the foundation denoising networks and generative facial prior. To demonstrate the effectiveness of the pretrained Stable Diffusion weights, we train our BFRffusion without them for initializing the U-Net denoiser. A significant drop is observed and it can’t generate realistic glasses as shown in Fig. 8(g), which shows that the generative ability from the pretrained Stable Diffusion is quite important for the restoration process. We employ a training strategy where we finetune the frozen diffusion model for 100K iterations, then unfreeze the decoder weights of the U-Net in Stable Diffusion and train the whole restoration model for 150K iterations. This training strategy allows our BFRffusion to achieve wonderful realness and fidelity while retaining the generation ability of Stable Diffusion. We freeze the whole U-Net all the time leading to a bad optimization of the restoration model and a slight performance drop. On the other hand, unfreezing U-Net results in the forgetting of prior knowledge of the pretrained Stable Diffusion and a noticeable performance drop. In the U-Net architecture of diffusion models, the encoder is responsible for gradually downsampling the input into high-dimensional feature representation and extracting rich semantic information, while the decoder gradually restores the spatial details of the image through upsampling operations and performs prediction. We use the proposed MFEM to extract features from the input low-quality images, so We choose to unfreeze the decoder rather than the encoder to achieve better restoration results. Table V(j) shows that the unfrozen decoder yields 0.70 dB PSNR gain over the unfrozen encoder.
Figure 9: Limitations of our BFRffusion. Our BFRffusion struggles to restore clear images when degradation is severe or watermarks are existent.
IV-D Limitations
Fig. 9 shows some limitations of our BFRffusion. When the degradation of real-world face images is severe, artifacts may appear in the images restored by our BFRffusion. The reason is that our restoration model is trained on the synthetic degraded data, which may not cover all degradation scenarios encountered in the real world. In the future, we plan to utilize higher-quality data to train the restoration model. Additionally, when watermarks appear in face images, our BFRffusion struggles to restore these watermarks while restoring clear faces, significantly affecting the overall visual appearance of the images. It is the existence of these limitations that motivates researchers to continuously explore more effective blind face restoration methods.
TABLE VI: Quantitative comparison on CelebA-Test and PFHQ-Test for blind face restoration trained on FFHQ and our PFHQ dataset. Bold indicates better performance.
Methods Training CelebA-Test PFHQ-Test
@@ IV-D Limitations
HiFaceGAN FFHQ 24.92 0.6195 24.48 0.5509
PFHQ 25.04 0.6547 24.60 0.5984
PSFRGAN FFHQ 24.71 0.6557 24.65 0.5784
PFHQ 25.19 0.6519 24.70 0.5923
BFRffusion FFHQ 26.20 0.6926 25.24 0.6193
PFHQ 26.10 0.6876 25.26 0.6210
TABLE VII: Quantitative comparison on three real-world datasets for blind face restoration trained on FFHQ and PFHQ dataset. Bold indicates better performance.
Methods Training LFW-Test CelebAdult-Test WIDER-Test
@@
HiFaceGAN FFHQ 64.50 104.02 125.56
PFHQ 48.00 118.88 43.92
PSFRGAN FFHQ 51.89 104.08 51.16
PFHQ 48.23 121.39 43.38
BFRffusion FFHQ 49.92 103.64 56.97
PFHQ 58.19 118.10 50.39
Figure 10: Visual comparison results on synthetic testing datasets for methods trained on the FFHQ dataset and PFHQ dataset.
Figure 11: Visual comparison results on three real-world testing datasets for methods trained on the FFHQ dataset and PFHQ dataset.
IV-E Utility Evaluation of the Built Privacy-preserving Face Dataset
To evaluate our privacy-preserving face dataset PFHQ, we train three representative blind face restoration methods: HiFaceGAN [76] (non-prior-based methods), PSFRGAN [7] (geometric prior-based methods), and BFRffusion (generative prior-based methods) on PFHQ dataset and compare the performance of these methods trained on FFHQ dataset [19]. We test these methods on CelebA-Test, PFHQ-Test, and three real-world testing datasets mentioned in Section IV-B. The quantitative comparisons on the synthetic dataset CelebA-Test and PFHQ-Test are reported in Table VI. Bold values indicate better performance. Specifically, On the CelebA-Test, HiFaceGAN [76] trained on our PFHQ dataset achieves better performance than on the FFHQ dataset. PSFRGAN and BFRffusion trained on our PFHQ dataset achieve comparable performance to these when trained on the FFHQ dataset. On the balanced testing dataset PFHQ-Test, the methods trained on the PFHQ dataset achieve better performance than on the FFHQ dataset. The quantitative comparisons on the public real-world testing datasets are reported in Table VII. Three blind face restoration methods trained on our PFHQ dataset also achieve comparable performance to those trained on the FFHQ dataset. The qualitative results are presented in Fig. 10 and 11. Due to the balanced characteristics of our PFHQ dataset, methods trained on it have better applicability to face images of various races, genders, and ages.
@@ V Conclusion and Future Work
In this work, we propose BFRffusion with delicately designed architecture to leverage amazing generative priors encapsulated in pretrained Stable Diffusion for blind face restoration. Our BFRffusion effectively restores realistic and faithful facial details and achieves state-of-the-art performance on both synthetic and real-world public testing datasets. Furthermore, we build a privacy-preserving paired face dataset called PFHQ with balanced race, gender, and age. Extensive experiments show that our PFHQ dataset can serve as an alternative to real face datasets for training blind face restoration methods.